A merchant has coffee worth $5 a pound that she wishes to mix with 70 pounds of coffe worth $8 a pound to get a mixture that can be sold for $7 a pound. How many pounds of the $5 coffee should be used?
Answer by nerdybill(7384)

A merchant has coffee worth $5 a pound that she wishes to mix with 70 pounds of coffee worth $8 a pound to get a mixture that can be sold for $7 a pound. How many pounds of the $5 coffee should be used?
.
Let x = amount (pounds) of $5 coffee added
then
5x + 8(70) = 7(70+x)
5x + 560 = 490+7x
560 = 490+2x
70 = 2x
35 pounds = x (answer)